OpenStudy (anonymous):

I submitted this and the prof told me the results in one of my compositions is wrong. (f ° g)(x) and (g ° f)(x) f(x)=√(x+1),g(x)=x+2 f(x+2)=√(x+2+1) =√(x+3) =x+3 g(√(x+1))=√(x+1)+2 =x+1+2 =x+3

OpenStudy (anonymous):

f o g (x) = f ( sqrt (x+1) ) = sqrt [ sqrt (x+1) + 1 ]

OpenStudy (anonymous):

i messed up

OpenStudy (anonymous):

f o g (x) = f ( sqrt (x+1) ) = sqrt [ sqrt (x+1) + 1 ]

OpenStudy (anonymous):

correction f o g (x) = f ( x+2 ) = sqrt [ (x+2)+1) = sqrt ( x + 3)

OpenStudy (anonymous):

g o f ( x) = g ( sqrt (x+1) ) = sqrt (x+1) + 2

OpenStudy (anonymous):

you cant simplify that further